# Bin Yang1,2, Alexandru Paul Condurache1,2

### v1

@@ Abstract
3D scene understanding is a critical yet challenging task in autonomous driving, primarily due to the irregularity and sparsity of LiDAR data, as well as the computational demands of processing large-scale point clouds. Recent methods leverage the range-view representation to improve processing efficiency. To mitigate the performance drop caused by information loss inherent to the ”many-to-one” problem, where multiple nearby 3D points are mapped to the same 2D grids and only the closest is retained, prior works tend to choose a higher azimuth resolution for range-view projection. However, this can bring the drawback of reducing the proportion of pixels that carry information and heavier computation within the network. We argue that it is not the optimal solution and show that, in contrast, decreasing the resolution is more advantageous in both efficiency and accuracy. In this work, we present a comprehensive re-design of the workflow for range-view-based LiDAR semantic segmentation. Our approach addresses data representation, augmentation, and post-processing methods for improvements. Through extensive experiments on two public datasets, we demonstrate that our pipeline significantly enhances the performance of various network architectures over their baselines, paving the way for more effective LiDAR-based perception in autonomous systems. The code will be released based on the acceptance.
@@ 1 Introduction
In general, those methods can be categorized based on LiDAR data representation into point-based [47, 26, 39], voxel-based [49, 14] and projection-based methods [46, 24, 5, 48]. Both point- and voxel-based approaches typically require substantial computational resources due to the need to process data through networks with numerous 3D convolutional layers, intensive feature pre-processing, and deep architectures involving multiple downsampling and upsampling operations. These requirements can result in slow inference speeds, limiting their suitability for real-time applications. In contrast, rasterizing point cloud into range-view images [7] is more advantageous in fast and scalable LiDAR perception, because it allows the use of 2D operators for efficient computation and facilitates the transfer of knowledge from camera images [1, 18].
@@
Nevertheless, learning from range-view representations can suffer from the performance drop caused by the ”many-to-one” conflict of adjacent points. To resolve the problem, most approaches maximize the azimuth resolution to make the range image more informative [24, 7, 5, 48, 13, 18]. We argue that this is suboptimal for the following reasons: 1) High overhead: increasing resolution significantly raises computational demand, limiting real-time performance. 2) Inefficient computation: the sparse nature of LiDAR data leads to many empty grid cells. Those unoccupied pixels introduce noise in the data and consume resources needlessly. 3) Limited informative gains: due to distortions caused by ego-motion and sensor-internal errors, points mapped to the same azimuthal location do not perfectly align with the laser beams. This misalignment implies that prioritizing both azimuth and elevation resolutions, rather than increasing image width alone, would yield a higher projection rate, as seen in our analysis in Fig. 2.
@@
Building on the aforementioned insights, we introduce FLARES, a novel training schema for range-view semantic segmentation of LiDAR point clouds. Rather than focusing on designing a new network architecture, FLARES targets to address the common limitations outlined above, hence, our proposal is generalizable to any range-view-based approaches. In essence, FLARES divides the point cloud into multiple sub-clouds, with only one projected onto a low-resolution image during training. During inference, all sub-clouds are projected and stacked along the batch dimension for processing. This approach increases the projection rate in both azimuth and elevation, thereby enriching informativeness of the range-view representation at lower cost.
Downscaling resolution, however, can exacerbate class imbalance in the dataset [7], potentially leading to overfitting during training. Another problem is decreasing 2D occupancy due to splitting the point cloud. To tackle these two incidental issues, we extend the pipeline with two additional data augmentation steps. Furthermore, we explore improvements in post-processing methods, a topic that has received minimal attention in previous works.
In summary, our contributions are as follows: 1) We introduce FLARES, a newly designed training schema for faster and more accurate LiDAR semantic segmentation. 2) We integrate two data augmentation techniques tailored to the new schema to enhance the network performance. 3) We propose a novel interpolation-based post-processing approach to resolve the ”many-to-one” problem more effectively. 4) We generalize this schema and all components across various range-view semantic segmentation networks, demonstrating superior performance over baselines on two different benchmarks.
@@ 2 Related Works
Point- and Voxel-based methods Some recent works [26, 47, 38] use raw point cloud data as direct network input, eliminating the need for post-processing after prediction. However, these methods often face high computational complexity and memory usage. To address these issues, Hu et al. [17] introduced sub-sampling and feature aggregation techniques for large-scale point clouds to reduce computational costs. Despite these efforts, performance degradation remains significant. Other works [49, 19] use 3D voxel grids as input, achieving point-based accuracy with reduced computational costs by utilizing sparse 3D convolutions [6]. Nonetheless, voxelization and de-voxelization steps continue to be time- and memory-intensive. Range-view-based methods To address inefficiencies, some prior works [36, 24, 37] convert the large-scale point cloud to panoramic range image through spherical projection and leverage image segmentation techniques for LiDAR data. SalsaNext [7] uses a Unet-like network with dilated convolutions to broaden receptive fields for more accurate segmentation, while Lite-HDSeg [28] introduces an efficient framework using a lite version of harmonic convolutions. Additionally, FIDNet [48] and CENet  [5] interpolate and concatenate multi-scale features with a minimal decoder for semantic prediction. These methods share the benefit of lightweight network design, significantly improving efficiency and enabling real-time applications. Nevertheless, they generally underperform 3D methods due to the “many-to-one” issue, where multiple points project to the same pixel. To offset the performance drop caused by the problem, some other recent works propose to use Vision Transformer (ViT) [9, 41, 35]. RangeViT [1] deploys standard ViT backbone as encoder, followed by a light-weight decoder for refining the coarse patch-wise ViT representations, while RangeFormer [18] utilizes a pyramid-wise ViT-encoder to extract multi-scale features from range images. ViTs offer higher model capacities and excel at capturing long-range dependencies by modeling global interactions between different regions, enhancing segmentation performance over traditional CNNs [8]. However, the quadratic computational complexity of self-attention mechanisms in ViTs introduces challenges in achieving an optimal balance between efficiency and accuracy. Training schema Highlighting inefficiencies stemming from the use of high-resolution range images, some methods have adopted compact networks, significantly reducing network capacity [48, 7, 5]. Unfortunately, high-resolution range images still demand substantial memory, which restricts scalability in terms of batch size and data throughput. Lowering the resolution exacerbates information loss, leading to inferior results. To address this, Kong et al. [18] proposed Scalable Training from Range view (STR), a strategy that divides range images into multiple sub-images from different perspectives to reduce memory consumption. Although STR lessens memory usage, it results in a slight drop in segmentation accuracy and only minor improvement in inference speed compared to the baseline. Augmentation Data augmentation plays a crucial role in helping models learn more generalized representations, thereby enhancing scalability. For example, Mix3D [25] introduced an out-of-context mixing strategy by fusing two scenes. Similarly, MaskRange [13] proposed a weighted paste-drop augmentation to manually balance the class frequencies, while RangeFormer [18] employed four consecutive range-wise operations to provide richer semantic and structural cues in the scene. In this work, we introduce two novel augmentation steps, including point-wise and range-wise fusion, specifically designed for the range images utilized in our new schema. Post-Processing Addressing the prevalent ”many-to-one” problem in range-view representations often necessitates a post-processing step to upsample 2D predictions, a critical yet underexplored area in prior research. For example, RangeViT [1] introduced a trainable 3D refiner using KPConv [32]. However, while it directly optimizes 3D semantics, the performance improvement is limited, and the approach adds significant computational overhead. Some methods rely on conventional unsupervised techniques to infer semantics for 3D points; for instance, Milioto et al. [24] proposed a KNN-based voting approach, and Zhao et al. [48] introduced Nearest Label Assignment (NLA), which assigns labels based on the closest labeled point in 3D space. Nevertheless, these unsupervised techniques often struggle with accurately predicting boundaries and distant points, with performance further declining as range-image resolution decreases. To overcome these limitations, we design a new post-processing method that better interpolates the predictions of unprojected points using neighborhood information.
3 Proposed Approach
Our method introduces critical optimizations across three areas: a distinct training and inference scheme for low-resolution range images, advanced augmentation techniques, and an effective post-processing approach for accurately mapping 2D predictions into 3D space.
3.1 Pre-processing
Range-view Projection A LiDAR point cloud consists of points captured during a single revolution, denoted as $P=\{p_{1},...,p_{n}\}$, where each measurement represents a 4D point including the cartesian coordinates $p_{i}=\{x_{i},y_{i},z_{i}\}$ and intensity $t_{i}$. By pre-defining the 2D resolution, assuming $W$ and $H$ as the image width and height, we can project the point cloud into a range image, where each row $v$ and column $u$ correspond to the elevation and azimuth angles of LiDAR points. The mathematical expression of the spherical projection model is as following:
$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arcsin(\frac{z}{d}))\end{bmatrix}$ (1)
Angular values $\Theta_{max}$ and $\Theta_{min}$ define the upper and lower bound of the LiDAR’s vertical field of views and the depth value is calculated by $d=\sqrt{x^{2}+y^{2}+z^{2}}$. Note that $H$ is typically determined by the number of LiDAR sensor beams, while W can be assigned with random value based on the requirements. Similar to the prior studies [5, 48, 7], we adopt a five-channel input representation ($x,y,z,t,d$). Data Representation Increasing azimuth resolution can help preserve more details from point clouds, but this comes at the expense of reduced efficiency. Additionally, as illustrated in Fig. 2, elevation resolution plays an equally important role in maximizing projection rates. This led us to rethink if a large image width is necessary to mitigate information loss. Inspired by STR [18], we propose an alternative solution: downsampling the point cloud into multiple equal-interval sub-clouds and projecting them into range images with a lower azimuth resolution. During training, we randomly select one range image. For inference, we stack all images along the batch dimension and process them in a single forward pass. We depict the workflow of FLARES in Fig. 3. To balance 3D validity and 2D occupancy, assuming $N_{max}$ is the maximum number of partitioning groups for the specific resolution of range images, it is determined by the rule that the average 2D occupancy must not fall below the high-resolution range image in standard mode ($\frac{1}{N_{max}}\sum^{N_{max}}_{i}Occ_{i}\geq Occ_{high}$). This new design offers three key advantages: 1) Enhanced projection rate by increasing both image height and width; 2) Reduced memory consumption, enabling deployment on smaller GPUs; 3) Preservation of the full field of view, maintaining contextual integrity despite downsampling.
Figure 3: Visual Illustration of FLARES: The full LiDAR point cloud is equally divided and grouped, each projected into a lower-resolution range image. During training, one image and its 2D label are sampled for optimization. For testing, stacked sub-cloud projections are processed simultaneously, and the outputs are fused into 3D predictions using an unsupervised method. We provide more details in the supplementary material.
@@ 3.2 Data Augmentation
Previous studies [5, 1, 48] have primarily used geometric transformations such as random flipping, translation, and rotation for point cloud augmentation. To further enrich semantic contexts and optimize for the low-resolution range image used in FLARES, we introduce two additional augmentation steps: 1) Weighted Paste-Drop+ Class imbalance is a common issue in LiDAR semantic segmentation benchmarks [2, 3], where certain classes are heavily underrepresented. This imbalance is compounded by information loss during point cloud-to-range image conversion, especially when downscaling azimuth resolution. Building upon the Weighted Paste Drop (WPD) approach introduced in MaskRange [13], which pastes pixels from rare classes and drops pixels from abundant classes, we present an enhanced version, WPD+. Unlike the original method, which performs geometric data augmentation identically on both sampled and current frames in 3D space before projection, our approach applies WPD directly in 3D space, which avoids the repeating computation of geometric transformations, and samples multiple frames to improve class balancing. Additionally, we use a small set of synthetic dataset generated in the Carla Simulator [10] to further augment rare classes that correspond to small and dynamic objects in the scene. Despite possible domain gaps between datasets, it yields notable accuracy improvements from our experimental results. 2) Multi-Cloud Fusion Given the inherent sparsity of LiDAR point clouds, range images often contain a considerable number of empty grids. Prior works have mitigated the issue by interpolating missing pixels from nearest neighbors [44] or by supplementing occupied pixels from other frames [18]. Coming down to the point cloud splitting in FLARES mode, the 2D occupancy in the single range image is decreased and it prompts the necessity for some solution. In our empirical study in Fig. 5(d), we found that it yields sub-optimal results when training directly on sub-cloud. To address the issue, we propose Multi-Cloud Fusion (MCF), a strategy to increase the 2D occupancy of sub-clouds alongside the setup of FLARES. Assuming the point cloud can be divided into a maximum of $N_{max}$ groups, we randomly pick a group number $N$in $\{1,2,...,N_{max}\}$ and split the full point cloud accordingly. After projecting them into $N$ range images, we randomly select one $R\in\mathbb{R}^{H\times W\times 5}$ as the training input. To enhance occupancy, all empty pixels in $R$ are filled using occupied pixels from remaining $N-1$ range images. This method maximizes the 2D occupancy in the range image of a sub-cloud while maintaining the structural consistency of the scene.
For further technical details of how input data is curated and augmented, please refer to the supplementary material.

### v2

# ΥFLARESΥ: Fast and Accurate LiDAR Multi-Range Semantic Segmentation

@@ Abstract
3D scene understanding is a critical yet challenging task in autonomous driving due to the irregularity and sparsity of LiDAR data, as well as the computational demands of processing large-scale point clouds. Recent methods leverage range-view representations to enhance efficiency, but they often adopt higher azimuth resolutions to mitigate information loss during spherical projection, where only the closest point is retained for each 2D grid. However, processing wide panoramic range-view images remains inefficient and may introduce additional distortions. Our empirical analysis shows that training with multiple range images, obtained from splitting the full point cloud, improves both segmentation accuracy and computational efficiency. However, this approach also poses new challenges of exacerbated class imbalance and increase in projection artifacts. To address these, we introduce FLARES, a novel training paradigm that incorporates two tailored data augmentation techniques and a specialized post-processing method designed for multi-range settings. Extensive experiments demonstrate that FLARES is highly generalizable across different architectures, yielding 2.1%–7.9% mIoU improvements on SemanticKITTI and 1.8%–3.9% mIoU on nuScenes, while delivering over 40% speed-up in inference. [^1] Project page: https://binyang97.github.io/FLARES
@@ 1 Introduction
In general, those methods can be categorized based on LiDAR data representation into point-based [64, 38, 53], voxel-based [66, 20] and projection-based methods [63, 34, 7, 65]. Both point and voxel-based approaches typically require substantial computational resources due to the need to process data through networks with numerous 3D convolutional layers, intensive feature pre-processing, and deep architectures involving multiple downsampling and upsampling operations. These requirements can result in slow inference speeds, limiting their suitability for real-time applications [31]. In contrast, rasterizing point cloud into range-view images [10] is more advantageous in fast and scalable LiDAR perception, because it allows the use of 2D operators for efficient computation and facilitates the transfer of knowledge from camera images [2, 25].
(a) Standard
(b) STR [25]
(c) FLARES
@@
Nevertheless, learning from range-view representations can suffer from a ”many-to-one” conflict among adjacent points. This issue causes an irreversible loss of spatial information and leads to inferior performance compared to other methods that utilize 3D LiDAR representations. Although increasing the azimuth resolution of range images can alleviate this loss, it forces the network to process wide panoramic images (Fig. 1(a)), which significantly increases computational overhead [34, 10, 7, 65, 19, 25]. Building on this standard setting, recent work introduces a scalable training framework with range-view images (STR) to improve efficiency [25]. As shown in Fig. 1(b), by splitting the point clouds into multiple distinct views, STR processes all crops in a batch during inference. The resulting predictions are then merged to reconstruct the full range image, which is subsequently post-processed to generate 3D coordinates.
However, the splitting strategy of STR is heuristic and compromises scene completeness. Empirical results in [25] reveal a slight performance drop within some networks. Furthermore, the gain in inference speed is limited when maintaining the high resolution of range images during the projection and post-processing stages. To solve these challenges, we propose FLARES, a novel training paradigm for range-view LiDAR semantic segmentation. Its procedure is visually illustrated in Fig. 1(c). In contrast to cartesian grid splits of STR [25], FLARES divides point clouds along the LiDAR’s spherical coordinate and then projects each split into lower-resolution range images. This design preserves local geometric coherence and scene integrity during partitioning, which enables more effective semantic context sharing across generated range images compared to STR. Furthermore, our method exclusively processes low-resolution images, achieving greater efficiency than both STR and standard settings.
@@
Moreover, multi low-resolution range-view images created by FLARES are generally a more informative and powerful representation compared to the high-resolution projection. Due to ego-motion and sensor calibration error, points mapped to identical azimuthal positions often misalign with actual laser beam orientations [61]. This inherent mismatch demonstrates that jointly prioritizing azimuth and elevation resolutions, rather than solely expanding image width, achieves higher projection fidelity. Statistically, as quantified in our occupancy measurement (Fig. 2), doubling the elevation resolution achieves approximately 80% of the projection gain obtained by doubling the azimuth resolution in SemanticKITTI [3] dataset. This property indicates that our new representation can restore more 3D points, thereby alleviating the ”many-to-one” conflict among adjacent points more effectively.
(a) Before splitting
(b) After splitting
Figure 3: An example from SemanticKITTI [3] dataset is visualized in top-down view and range-view. Three crops are magnified to specify three limitations brought by the splitting of point clouds. 1) Exacerbated class imbalance: objects with small sizes, likewise low occurrences in annotations, tend to fade away in the range image after downsampling (e.g. pole in the left crop). 2) Intensified noise: reduction in points results in increasing clutters (middle crop) that may disrupt the training stability. 3) Deteriorated distortion: decrease in point density can introduce more projection artifacts, thus corrupting the sharpness of local geometry (e.g. blurry boundaries of car in the right crop).
While FLARES offers significant advantages, it introduces three key challenges, as visualized in Fig. 3. First, point cloud splitting exacerbates class imbalance issues [6], where classes representing small objects with inherently sparse distributions become further undersampled. Second, the splitting operation increases point cloud sparsity, causing substantial reductions in occupancy of range images. These unoccupied pixels act as noise that destabilize training convergence [25]. Finally, the resulting projection artifacts degrade image sharpness, particularly at partition boundaries where geometric discontinuities emerge.
To tackle these incidental issues, we extend the pipeline with two data augmentation steps. Furthermore, we investigate in exploring a more effective post-processing technique tailored to the multi-range settings to further enhance the advantages of our method.
In summary, our contributions are:
  1. 1.
    FLARES, a scalable training paradigm that enhances LiDAR semantic segmentation with improved speed, accuracy, and architectural agnosticism.
  2. 2.
    Two specialized data augmentation techniques addressing exacerbated class imbalance and noise amplification within FLARES.
  3. 3.
    A novel interpolation-based post-processing method that refines range-view predictions using precise nearest-neighbor information.
  4. 4.
    Comprehensive experiments across four network architectures and two benchmarks demonstrating consistent superiority over baselines.
@@ 2 Related Works
Point and Voxel-based methods Some recent works [38, 64, 52] use raw point cloud data as direct network input, eliminating the need for post-processing after prediction. However, these methods often face high computational complexity and memory usage. To address these issues, Hu et al. [23] introduced sub-sampling and feature aggregation techniques for large-scale point clouds to reduce computational costs. Despite these efforts, performance degradation remains significant. Other works [66, 26] use 3D voxel grids as input, achieving point-based accuracy with reduced computational costs by utilizing sparse 3D convolutions [8]. Nonetheless, voxelization and de-voxelization steps continue to be time- and memory-intensive.
@@
Range-view-based methods To address inefficiencies, some prior works [50, 34, 51] convert the large-scale point cloud to panoramic range image through spherical projection and leverage image segmentation techniques for LiDAR data. SalsaNext [10] uses a Unet-like network with dilated convolutions to broaden receptive fields for more accurate segmentation, while Lite-HDSeg [41] introduces an efficient framework using a lite version of harmonic convolutions. Additionally, FIDNet [65] and CENet  [7] interpolate and concatenate multi-scale features with a minimal decoder for semantic prediction. These methods share the benefit of lightweight network design, significantly improving efficiency and enabling real-time applications. Nevertheless, they generally underperform 3D methods due to the “many-to-one” issue, where multiple points project to the same pixel. To offset the performance drop caused by the problem, some other recent works propose to use Vision Transformer (ViT) [13, 55, 49]. RangeViT [2] deploys standard ViT backbone as encoder, followed by a light-weight decoder for refining the coarse patch-wise ViT representations, while RangeFormer [25] utilizes a pyramid-wise ViT-encoder to extract multi-scale features from range images. ViTs offer higher model capacities and excel at capturing long-range dependencies by modeling global interactions between different regions, enhancing segmentation performance over traditional CNNs [11]. However, the quadratic computational complexity of self-attention mechanisms in ViTs introduces challenges in achieving an optimal balance between efficiency and accuracy.
Training Paradigm While existing methods address inefficiencies in high-resolution range images by employing compact networks to reduce model capacity [65, 10, 7], these approaches still require substantial memory for processing high-resolution inputs, limiting scalability in batch size and data throughput. Reducing resolution mitigates memory demands but worsens spatial information loss, degrading segmentation accuracy. To resolve this trade-off, Kong et al. [25] introduced Scalable Training from Range-view (STR), which splits range images into sub-images from multiple perspectives to lower memory consumption. However, STR’s use of partial scene views during training inherently limits segmentation accuracy. Furthermore, the method does not optimize for computational efficiency, leaving runtime improvements unaddressed.
Post-Processing Addressing the prevalent ”many-to-one” problem in range-view representations often necessitates a post-processing step to upsample 2D predictions, a critical yet under-explored area in prior research. RangeViT [2] introduced a trainable 3D refiner using KPConv [46]. However, while it directly optimizes 3D semantics, the performance improvement is limited, and the approach adds significant computational overhead. Some methods rely on conventional unsupervised techniques to infer semantics for 3D points. For instance, Milioto et al. [34] proposed a KNN-based voting approach, and Zhao et al. [65] introduced Nearest Label Assignment (NLA), which assigns labels based on the closest labeled point in 3D space. Nevertheless, these unsupervised techniques often struggle with accurately predicting boundaries and distant points, with performance further declining as range-image resolution decreases. To overcome these limitations, we design a new post-processing method that adaptively weights neighboring contributions for interpolating predictions in 3D space.
3 Methodology
3.1 Proposed Framework
A LiDAR point cloud consists of points captured during a single revolution, denoted as $P=\{p_{1},...,p_{n}\}$, where each measurement represents a 4D point including the cartesian coordinates $p_{i}=\{x_{i},y_{i},z_{i}\}$ and intensity $t_{i}$. In semantic segmentation task, each point has an additional feature of class label, denoted as $c_{i}$, for training. Our splitting configuration operates as follows: given a predefined partition count $N$, we distribute points into sub-clouds by along the spherical coordinate of the scan pattern. Specifically, a sub-cloud is derived as $P_{i}=\{p_{j}|j\mod N=i-1\}$ for $i\in\{1,2,..N\}$. This modulo-based assignment ensures equal partitioning while preserving spatial coherence within each sub-cloud. For the spherical projection, we specify the function $f$ as following:
$f(p)=\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arcsin(\frac{z}{d}))\end{bmatrix}$ (1)
, where $W$ and $H$ as the image width and height, while $v$ and $u$ correspond to the elevation and azimuth angles of LiDAR points. Angular values $\Theta_{max}$ and $\Theta_{min}$ define the upper and lower bound of the LiDAR’s vertical field of views and the depth value is calculated by $d=\sqrt{x^{2}+y^{2}+z^{2}}$. Note that $H$ is typically determined by the number of LiDAR sensor beams, while $W$ can be assigned with random value based on the requirements. Similar to prior studies [7, 65, 10], we adopt a five-channel input representation ($x,y,z,t,d$).
Next, we project all sub-clouds into range images simultaneously $R_{i}=f(P_{i})$. During training, since our splitting strategy leverages shared semantic contexts across sub-clouds, we randomly select only one for network optimization. During inference, all images $R_{i}$ are stacked as a batch and processed by the network to extract the 2D predictions. Unlike STR [25], which wraps all images to the original resolution, we treat each sub-cloud as an independent complete scene and apply post-processing to low-resolution range images individually. Final 3D predictions are reconstructed by merging outputs from all sub-clouds.
An discussed in Sec. 1, FLARES offers three key advantages: 1) Enhanced restoration of original 3D information through multi-range projection, 2) Reduced memory consumption by decreasing the image resolution, enabling the scalable training on smaller GPUs, 3) Preservation of the full field of view, which maintains contextual integrity despite downsampling.
@@ 3.2 Data Augmentation
In particular, the point cloud splitting strategy of FLARES implicitly introduces two new challenges: exacerbated class imbalance and amplified projection noise. To address class imbalance, prior works [19, 15, 57] have primarily proposed augmenting long-tail classes by pasting them from other frames. While effective to some extent, we find that the augmentation strength is insufficient to counteract the heightened imbalance in our setting. To this end, we propose an enhanced data augmentation strategy specifically designed for class balancing. As for projection noise, recent approaches [58] attempt to mitigate it by interpolating pseudo pixels into unoccupied regions. However, this often introduces new artifacts in the range-view image. In contrast, tailored to the multi-range configuration of FLARES, we leverage the unused sub-clouds to increase pixel occupancy in the sampled range-view image, thereby reducing projection artifacts.
Weighted Paste-Drop+ (WPD+) To address the problem of class imbalance, a pervasive challenge in LiDAR semantic segmentation benchmarks [3, 16], Weighted Paste-Drop (WPD) [19] technique provides a simple yet effective solution by selectively pasting pixels from rare classes while dropping pixels from abundant ones. Building on this approach, we introduce WPD+, a fundamentally re-engineered augmentation strategy specifically designed for FLARES.
Unlike traditional methods that apply augmentation uniformly on range-view images [19, 25], WPD+ operates directly in 3D space, This design not only avoids repeated geometric transformations on sampled and current frames, but also harnesses multiple scenes to more effectively balance class distributions and mitigate the severe imbalance induced by point cloud splitting.
To further enhance occurrence of underrepresented classes, particularly those corresponding to small and dynamic objects in the scene, we incorporate a curated synthetic dataset from the Carla Simulator [14] (more details are provided in Appendix A.5). From our empirical results, despite inherent domain gaps, this approach has led to notable accuracy gains in long-tail classes.
Multi-Cloud Fusion (MCF) Due to the point cloud splitting in FLARES, the decrease in 2D occupancy leads to intensified projection artifacts in range images. To address the issue, we propose MCF, a strategy tailored to multi-range settings of FLARES. We formalize this process as follows. Let $R_{i}:\Omega\to\mathbb{R}\cup\{\varnothing\}$ denote the range image from the $i$-th sub-cloud over the pixel domain $\Omega$. For each pixel $x\in\Omega$, We first define the set using other range images:
$S(x)=\{\,R_{j}(x)\mid j\neq i\text{ and }R_{j}(x)\neq\varnothing\,\}.$
Then, for a range image $R_{i}$ over the pixel domain $\Omega$, the occupancy-filled value $\tilde{R}_{i}(x)$ is given by:
$\tilde{R}_{i}(x)=\begin{cases}R_{i}(x),&\text{if }R_{i}(x)\neq\varnothing,\\[4.30554pt] f(S(x)),&\text{if }R_{i}(x)=\varnothing\text{ and }S(x)\neq\emptyset,\\[4.30554pt] \varnothing,&\text{otherwise.}\end{cases}$
Here, $f(\cdot)$ is an aggregation function that combines the occupied pixel values from the remaining N1N-1 range images. Consequently, this method maximizes the 2D occupancy in the range image of a sub-cloud while maintaining the structural consistency of the scene.
Despite their simplicity, both augmentation strategies are effective in enhancing segmentation performance. Moreover, when combined with the multi-range representations of FLARES, they significantly outperform prior methods (see Appendix B.3 for related experiments).